OpenStudy (anonymous):

Liz is a hairstylist. Overall, 36% of her clients get their hair colored. If 18% of her clients get their hair colored, 48% of her female clients get their hair colored, what percent of her clients are male?

OpenStudy (mathstudent55):

36% of her clients get hair colored 18% of clients get hair colored. Are these male of female?

OpenStudy (anonymous):

sorry 18% of the male clients get their hair colored

OpenStudy (mathstudent55):

Ok. M = number of male clients; F = number of female clients; T = total number of clients. M + F = T 36%T = 18%M + 48%F 0.36(M + F) = 0.18M + 0.48F 0.36M + 0.36F = 0.18M + 0.48F 0.18M = 0.12F M/F = 0.12/0.18 = 2/3 M/F = 2/3 Since the male to female ratio is 2 to 3, the male to total client ratio is 2 to (2 + 3), so it's 2/5. A 2/5 ratio is 40%